# FOCUS: Fused Observation of Channels for Unveiling Spectra

### v1

@@ 4.2 Main Results
As shown in Table 2, FOCUS incurs less than 1% parameter overhead and achieves the lowest FLOPs and runtime latency among all baselines. Importantly, it operates in a fully gradient-free manner at inference, which sharply reduces deployment cost compared to gradient-based methods (e.g., Grad-CAM [32], SANet [13]). These findings are consistent with recent work [8, 18], highlighting that prompt-based and gradient-free explainers enable practical, real-time visual interpretation for large transformer models.
4.3 Ablation Study
Table 3: Spectral prompt granularity. Moderate group count ($G=10$) achieves best interpretability without excess overhead.
#Groups $G$ Acc $\uparrow$ BIO@5 $\uparrow$ $\rho$uparrow Params (%)
@@ 4.3 Ablation Study
Table 4: Sink token and loss. Both accuracy and explanation stability improve with explicit sink routing.
Config Acc $\uparrow$ BIO@5 $\uparrow$ Sink Rate $\downarrow$ Sink Consist. $\uparrow$
w/o Sink 93.5 0.51 0.38 0.64
+ Sink, no loss 93.4 0.53 0.31 0.69
+ Sink & Loss 93.7 0.56 0.22 0.75
Spectral Prompt Granularity.
Table 3 shows that increasing the number of spectral groups ($G$) improves both accuracy and interpretability, but only up to a moderate value ($G=10$), beyond which gains saturate or slightly decline. This echoes prior findings [18, 13] that over-partitioning can dilute spectral focus and increase model overhead.
Sink Control Mechanism.
Table 4 demonstrates that introducing a dedicated [SINK] token and the associated attraction loss yields clear improvements in both interpretability and explanation stability, reducing spurious attention accumulation. This is consistent with recent work on attention regularization and prompt-based filtering in transformers [40, 10, 18].
4.4 Qualitative Analysis
We present visual comparisons to qualitatively evaluate the interpretability and spectral coherence of FOCUS. FOCUS produces cleaner and more localized saliency patterns than existing methods. Figure 3 illustrates that Grad-CAM tends to generate weak, spatially diffuse saliency, offering limited interpretive value under challenging spectral conditions. Prompt-CAM shows improved contrast but still lacks fine-grained localization in biologically informative regions.
In contrast, FOCUS produces sharp, consistent attention maps that tightly align with known physiological markers—specifically, stress-induced lesions in the spatial domain and well-defined spectral responses at the red-edge (750 nm) and SWIR ( 2,100 nm) bands, which are strongly linked to chlorophyll degradation and protein absorption in plant tissues. This highlights the biological validity and interpretability of our approach. In Figure 4, we observe that FOCUS is robust across replicates: attention consistently focuses on biochemically informative areas, whereas Prompt-CAM shows unstable activation and Grad-CAM fails to distinguish background noise. This suggests that our spectral prompts and sink routing modules provide more stable and reproducible attention allocation. Finally, Figure 6 shows band-wise heatmaps under different spectral groupings. FOCUS not only identifies where important features lie spatially, but also which specific bands matter—e.g., sharp peaks in the red-edge and protein-sensitive SWIR bands—demonstrating fine-grained spatial–spectral attribution. By contrast, baseline methods either collapse saliency across bands or highlight spectrally irrelevant regions, offering little insight to domain experts. Together, these qualitative findings reinforce the quantitative results and validate FOCUS as a reliable tool for interpretable hyperspectral vision tasks.
4.5 Analysis of Attention Sink Behavior
To elucidate how attention collapse manifests in ViT models on hyperspectral data, we analyze the distribution of attention mass routed to the [CLS] and [SINK] tokens across layers and heads. Figure 7 visualizes the per-head attention entropy and sink absorption rate under three configurations: (i) vanilla ViT, (ii) ViT with [SINK], and (iii) ViT with both [SINK] and the proposed attraction loss. Without explicit sink routing, over 30% of attention mass consistently accumulates on the first token, echoing the “attention-sink” phenomenon observed in recent ViT literature [40, 48]. By introducing our structured sink mechanism, this rate drops sharply, while predictive performance remains unchanged (see Table 4). This demonstrates that FOCUS does not merely regularize attention, but actively enables stable, trustworthy, and interpretable attention routing for long-sequence ViTs. Similar findings have been reported in recent XAI studies that combine architectural innovations with explicit routing or gating for robust interpretability [8, 18].
4.6 Robustness Under Spectral Noise
Hyperspectral imaging is inherently susceptible to noise at high wavelengths (especially SWIR $>$2200 nm) due to sensor artifacts, water absorption, or atmospheric effects. This not only reduces model reliability, but can also destabilize saliency explanations as transformers overfit to spurious spectral patterns. Empirically, we observe that vanilla ViT-based saliency methods (e.g., Grad-CAM [32], Prompt-CAM [8]) often produce volatile or physiologically irrelevant attributions in these noisy bands—a phenomenon also documented in [48, 5].
@@ 4.6 Robustness Under Spectral Noise
In contrast, our sink mechanism acts as a noise valve: unstable heads route high-entropy responses to [SINK], selectively suppressing noise-induced activations. This yields more robust and biologically meaningful explanations, without additional model tuning. Our results resonate with the “attention debiasing” literature, which finds that targeted attention steering—whether by loss design [5] or prompt gating [2]—improves model resilience to common corruptions. Uniquely, FOCUS achieves these gains with less than 1% additional parameters and without retraining the ViT backbone.
4.7 Computational Efficiency Comparison
@@
Interpretability in hyperspectral vision must be efficient to support deployment in real-world agricultural and biological scenarios with limited resources. However, standard ViTs scale poorly for hyperspectral cubes: a single image may contain 200+ spectral bands, leading to quadratic growth in memory and attention cost. Direct inference with a vanilla ViT requires over 14,GB of GPU memory. Recent approaches like D-CHAG [36] use spectral aggregation to reduce computation, yet still consume 8.4,GB and require model modifications and retraining. In contrast, FOCUS introduces no backbone changes or gradients. Through lightweight spectral prompts and a learnable [SINK] token, it requires only 1.9,GB of memory and operates in a single forward pass. As shown in Figure 5, FOCUS reduces memory cost by over 85% while delivering fine-grained spatial–spectral explanations. This efficiency makes our method the first practically deployable interpretability solution for ViTs in high-dimensional hyperspectral vision.
@@
Figure 5: Inference efficiency comparison. FOCUS achieves over 95% runtime and 85% memory reduction compared to vanilla ViT and D-CHAG [36], while providing complete spatial–spectral interpretability.
4.8 Limitations
Despite its strong interpretability and low overhead, FOCUS has two current limitations. First, the number of spectral prompts grows linearly with the number of target classes ($K$), which may impact scalability in large-scale recognition tasks. Future work may explore shared or hierarchical prompting strategies to reduce prompt complexity. Second, under severe spectral noise (e.g., beyond 2200 nm in SWIR), the [SINK] token may over-filter faint yet informative signals. Integrating adaptive routing or uncertainty-aware mechanisms could offer better trade-offs between suppression and retention.
@@ 4.8 Limitations
Figure 7: Layer-wise sink attention mass. Comparison of attention mass routed to the [SINK] token across transformer layers. Our full method (+Sink+$\mathcal{L}_{\text{sink}}$) most effectively drains non-discriminative attention, reducing accumulation over depth.
5 Conclusion
We introduced FOCUS, a lightweight and gradient-free interpretability framework tailored for hyperspectral vision transformers. By combining class-specific spectral prompts with a learnable [SINK] token and a simple attraction loss, FOCUS produces stable and fine-grained spatial–spectral saliency maps. Despite adding less than 1% additional parameters, it achieves state-of-the-art interpretability performance—improving spectral IoU by 15% and enhancing attribution consistency across layers. Our work addresses a fundamental scalability barrier in hyperspectral interpretability, enabling trustworthy and deployable explanations for high-dimensional spectral data. We believe FOCUS lays a practical foundation for interpretable AI in spectral-intensive domains. In future work, we aim to explore adaptive prompt sharing and hierarchical strategies to further scale FOCUS to long-sequence spectral tasks.
@@ Acknowledgments
This manuscript was authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.

### v2

@@ III-B Main Results
As shown in Table II, FOCUS incurs less than 1% parameter overhead and achieves the lowest FLOPs and runtime latency among all baselines. Importantly, it operates in a fully gradient-free manner at inference, which sharply reduces deployment cost compared to gradient-based methods (e.g., Grad-CAM [1], SANet [11]). These findings are consistent with recent work [3, 8], highlighting that prompt-based and gradient-free explainers enable practical, real-time visual interpretation for large transformer models.
III-C Ablation Study
Spectral Prompt Granularity.
Table IV shows that increasing the number of spectral groups ($G$) improves both accuracy and interpretability, but only up to a moderate value ($G=10$), beyond which gains saturate or slightly decline. This echoes prior findings [8, 11] that over-partitioning can dilute spectral focus and increase model overhead.
Sink Control Mechanism.
Table III demonstrates that introducing a dedicated [SINK] token and the associated attraction loss yields clear improvements in both interpretability and explanation stability, reducing spurious attention accumulation. This is consistent with recent work on attention regularization and prompt-based filtering in transformers [5, 13, 8].
III-D Qualitative Analysis
We use visual comparisons to evaluate the interpretability and spectral coherence of FOCUS. Overall, FOCUS produces cleaner and more localized saliency patterns than existing methods. As shown in Figure 3, Grad-CAM tends to generate diffuse saliency maps that offer limited value under challenging spectral conditions. Prompt-CAM improves contrast but still struggles to achieve fine-grained localization in biologically informative regions. In contrast, FOCUS generates sharp and consistent attention maps. These maps align closely with known physiological markers. Specifically, the model accurately highlights stress-induced lesions in the spatial domain. It also identifies distinct spectral responses in the red-edge (750 nm) and SWIR ($\sim$2,100 nm) bands. These wavelengths are linked to chlorophyll degradation and protein absorption in plant tissues, which supports the biological validity of our approach. Figure 4 demonstrates the robustness of FOCUS across replicates. The attention consistently targets biochemically informative areas. Conversely, Prompt-CAM exhibits unstable activation, and Grad-CAM fails to distinguish the object from background noise. This indicates that our spectral prompts and sink routing mechanism ensure more stable attention allocation. Finally, Figure 6 displays band-wise heatmaps for different spectral groups. FOCUS identifies not only where important features lie spatially but also which specific bands are significant. For example, it reveals sharp peaks in the red-edge and protein-sensitive SWIR bands. In comparison, baseline methods often collapse saliency across bands or highlight spectrally irrelevant regions, which provides limited insight for domain experts. These qualitative findings reinforce our quantitative results and validate FOCUS as a reliable tool for interpretable hyperspectral vision.
TABLE III: Sink token and loss. Accuracy and explanation stability improve with explicit sink routing.
Config Acc $\uparrow$ BIO@5 $\uparrow$ Sink Rate $\downarrow$ Consist. $\uparrow$
w/o Sink 93.5 0.51 0.38 0.64
+ Sink, no loss 93.4 0.53 0.31 0.69
+ Sink & Loss 93.7 0.56 0.22 0.75
TABLE IV: Spectral prompt granularity. G=10G{=}10 achieves best interpretability without excess overhead.
$G$ Acc $\uparrow$ BIO@5 $\uparrow$ $\rho$ $\uparrow$ Params (%)
@@ III-D Qualitative Analysis
III-E Robustness Under Spectral Noise
@@
Hyperspectral imaging is inherently susceptible to noise at high wavelengths (especially SWIR >>2200 nm) due to sensor artifacts, water absorption, or atmospheric effects. This not only reduces model reliability, but can also destabilize saliency explanations as transformers overfit to spurious spectral patterns. Empirically, we observe that vanilla ViT-based saliency methods (e.g., Grad-CAM [1], Prompt-CAM [3]) often produce volatile or physiologically irrelevant attributions in these noisy bands—a phenomenon also documented in [14, 15]. In contrast, our sink mechanism acts as a noise valve: unstable heads route high-entropy responses to [SINK], selectively suppressing noise-induced activations. This yields more robust and biologically meaningful explanations, without additional model tuning. Our results resonate with the “attention debiasing” literature, which finds that targeted attention steering—whether by loss design [15] or prompt gating [16]—improves model resilience to common corruptions. Uniquely, FOCUS achieves these gains with less than 1% additional parameters and without retraining the ViT backbone.
III-F Computational Efficiency Comparison
@@
Interpretability in hyperspectral vision must be efficient to support deployment in real-world agricultural and biological scenarios with limited resources. However, standard ViTs scale poorly for hyperspectral cubes: a single image may contain 200+ spectral bands, leading to quadratic growth in memory and attention cost. Direct inference with a vanilla ViT requires over 14,GB of GPU memory. Recent approaches like D-CHAG [17] use spectral aggregation to reduce computation, yet still consume 8.4,GB and require model modifications and retraining. In contrast, FOCUS introduces no backbone changes or gradients. Through lightweight spectral prompts and a learnable [SINK] token, it requires only 1.9,GB of memory and operates in a single forward pass. As shown in Figure 5, FOCUS reduces memory cost by over 85% while delivering fine-grained spatial spectral explanations. This efficiency makes our method the first practically deployable interpretability solution for ViTs in high-dimensional hyperspectral vision.
@@
Figure 5: Inference efficiency comparison. FOCUS achieves over 95% runtime and 85% memory reduction compared to vanilla ViT and D-CHAG [17], while providing complete spatial spectral interpretability.
@@
IV Conclusion
We introduced FOCUS, a lightweight and gradient-free interpretability framework tailored for hyperspectral vision transformers. By combining class-specific spectral prompts with a learnable [SINK] token and a simple attraction loss, FOCUS produces stable and fine-grained spatial–spectral saliency maps. Despite adding less than 1% additional parameters, it achieves state-of-the-art interpretability performance—improving spectral IoU by 15% and enhancing attribution consistency across layers. Our work addresses a fundamental scalability barrier in hyperspectral interpretability, enabling trustworthy and deployable explanations for high-dimensional spectral data. We believe FOCUS lays a practical foundation for interpretable AI in spectral-intensive domains. In future work, we aim to explore adaptive prompt sharing and hierarchical strategies to further scale FOCUS to long-sequence spectral tasks.In the future, we can extend our work into reinforcement learning or multi-agent systems to further enhance the interpretability of hyperspectral images [18, 19].
@@ Acknowledgments
This manuscript was co-authored by Oak Ridge National Laboratory (ORNL), operated by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. Any subjective views or opinions expressed in this paper do not necessarily represent those of the U.S. Department of Energy or the United States Government.